OpenStudy (anonymous):

simplify: -(8x+4)+5x

OpenStudy (skullpatrol):

any ideas?

OpenStudy (anonymous):

same as last time right?

OpenStudy (elise_a18):

Do you know how to combine like terms?

OpenStudy (anonymous):

−(8x+4)+5x distribute =−8x+−4+5x combine Like Terms =−8x+−4+5x =(−8x+5x)+(−4) =−3x+−4 answer =−3x−4

OpenStudy (skullpatrol):

$$\Huge -(8x+4)+5x=-1(8x+4)+5x$$

OpenStudy (elise_a18):

okay then...

OpenStudy (skullpatrol):

distribute the -1